OpenStudy (anonymous):

why is x-0.4x=96 0.6x=96

OpenStudy (anonymous):

\[x-0.4x==96\]\[1\times x - 0.4 \times x=96\]\[(1-0.4)x=96\]\[0.6x=96\]

OpenStudy (akashdeepdeb):

@amirreza1870 explained it very well. We use the reverse of the distributive property. x can be written as 1 * x. Grouping together x - 0.4x as (1-0.4)x gives us 0.6x. :)

OpenStudy (anonymous):

Got it!! Thank you, that was driving me nuts!!

OpenStudy (akashdeepdeb):

:)